Lewis Company’s standard labor cost of producing one unit of Product DD is 3.20 hours at the rate of $12.10 per hour. During August, 44,000 hours of labor are incurred at a cost of $12.30 per hour to produce 13,500 units of Product DD.

(a) Compute the total labor variance.

Total labor variance $enter the Total labor variance in dollars select an option

(b) Compute the labor price and quantity variances.

Labor price variance $enter a dollar amount select an option

Labor quantity variance $enter a dollar amount select an option

(c) Compute the labor price and quantity variances, assuming the standard is 3.40 hours of direct labor at $12.45 per hour.

Labor price variance $enter a dollar amount select an option

Labor quantity variance $enter a dollar amount select an option
Answer #1

a) Total labor variance = (13500*3.2*12.10)-(44000*12.30) = 18480 U

b) Labor price variance = (12.10-12.30)*44000 = 8800 U

Labor quantity variance = (43200-44000)*12.10 = 9680 U

c) Labor price variance = (12.45-12.30)*44000 = 6600 F

Labor quantity variance = (13500*3.4-44000)*12.45 = 23655 F